# Detecting environmental effects in gravitational waves from binaries perturbed by periodic forces

### v2

@@ IV.1.1 Comparison with secular dephasing
The previous equation speaks to the high potential detectability of secular dephasing. It suggests that, indeed, cumulative shifts of order few $1/\text{SNR}$ may be detectable. However, this conclusion ignores the fact that secular dephasing, by its own nature of being slowly accumulated, can be degenerate with many binary parameters that act in a similar way. In particular, it is highly degenerate with the binary’s initial phase, along with any post-Newtonian modifications to the binary’s chirp behaviour with a similar magnitude or frequency scaling. A more conservative requirement for the detectability of secular dephasing is that it should exceed a full cycle, i.e. $2\pi$, while many works have also highlighted cases where several cycles of dephasing are necessary to perform parameter estimation [see, e.g., 124, 72]. Here we use $2\pi$ as a detectability criterion, finding that the threshold acceleration for secular de-phasing amounts to:
@@
where $N_{\rm{c}}$ is the number of cycles at the frequency $f$. Interestingly, this suggests that periodic perturbations may be the dominant EE for high SNR sources that complete comparatively fewer cycles in the sensitivity band of their respective detectors. As we will see in Sec. V, this may be realised in several astrophysical scenarios. The euation above also suggests that, in sources where both secular and periodic forces are present, the latter must be significantly higher in amplitude to be detected at the same $\delta$SNR. However a more precise statement on detectability requires a careful Bayesian analysis. Indeed, as we will see in IV.4, these naive expectations are often misleading.
@@ IV.2 Detectability for chirping LISA sources
When stepping beyond the monochromatic case, it is necessary to account for the full noise power spectral density of a detector. Here we produce some simple detectability criteria for the recently adopted, space based gravitational wave detector LISA [4, 33], but note that similar estimates apply to any similar milli-Hz GW instrument [e.g. TianQin 87].
@@
$\sim 8$ $\implies\text{1$\sigma$ uncertainty < 0.5 for $\log_{10}(B^{\prime})$ and $\beta^{\prime}$}$
@@ IV.4 Test case: Degeneracies in secular vs periodic dephasing
The test case showcased in Fig. 6 is representative of all $\delta$SNR larger than a few, where the widths of the Gaussian posteriors roughly scale as $1/\delta$SNR. This suggests that the additional parameters required to describe periodic de-phasing are completely orthogonal to the usual vacuum binary parameters. Crucially, they do not cause any visible degradation in the recovery of the vacuum waveform.
@@
We are aware that a complete validation of this claim would require a more systematic study of both secular and periodic dephasing, for many sources and across an array of different waveform models, which is beyond the scope of this work. Nevertheless, the results above do suggest that waveforms of the type presented in Eqs. 27 and 28 represent a strong candidate to search for EE in GW signals without necessarily degrading the estimation of the vacuum parameters of the source.
Figure 7: Full showcase of the marginalised posteriors when attempting to recover a periodic (black) and a secular (red) dephasing. The injected periodic dephasing is characterised by $B^{\prime}=0.1$ and $\beta$ has a $\delta$SNR of approximately 40. the injected secular dephasing has the same amplitude and scaling, resulting in a $\delta$SNR of approximately 60. Nevertheless, the latter cannot be reconstructed confidently as it suffers from a degeneracy with the binary’s initial phase and spin parameter. The two MCMCs are performed with 32 parallel walkers for 50’000 steps, requiring approximately 90 minutes of compute on a single core. See Eqs. 45 for the meaning of the nine parameters.
@@ V.1 Generic multipole perturbations
We apply the formalism developed in this work to a very simple, heuristic description of multi-polar modifications to the gravitational potential. We imagine the vacuum gravitational field to be perturbed by a series of additional components, arising either from the presence of background fluid, field or from modifications to GR. As the binary orbits in the modified potential, it will experience time varying forces with the following physical scaling:
@@
Extrapolating from the monochromatic results discussed in section IV.1 we conclude that by employing waveforms of the type showcased by Eqs. 27 and 28, one should expect to be able to constrain completely general multipolar modifications to the gravitational potential (and therefore the spacetime metric), provided they exceed a certain typical magnitude:
@@ V.2 Dynamical and stellar triple systems
Triple systems of compact objects are a rich laboratory of gravitational dynamics, and are being studied in great detail in the context of the dynamical formation channel for GW sources [98, 17, 116, 82, 104, 105, 105, 8, 94, 7, 122, 111, 112, 134, 103, 81, 67, 113, 141, 139, 66, 21, 139]. It is known that a close-by tertiary massive body may induce several detectable imprints on GW emission from the inner binary. In this context, the most studied EEs are e.g. Doppler modulations that arise as a consequence of Römer delays or line-of-sight accelerations, as well as the secular influence of tidal fields throughout the inspiral [153, 20, 109, 69, 102, 27, 100, 147, 128, 50, 137, 126, 150, 76, 133, 115, 115]. Here we neglect those effects, focusing instead on the instantaneous forces experienced by the inner binary components, as they orbit in the time-varying gravitational potential induced by the tertiary.
In the top panel of Fig. 8 we show the time series of the instantaneous torque experienced by an inner binary due to its tertiary companion. In this specific case, the inner binary has masses of 25 and 20 $M_{\odot}$, while the outer companion has a mass of 30 $M_{\odot}$. The inner binary has a semi-major axis of $2.5\times 10^{-4}$ AU, and emits GWs at 0.1 Hz. The outer binary has a semi-major axis of $1.5\times 10^{-3}$, yielding a period ratio of approximately 11. The simulations are performed with the regularized, direct $N$-body code tsunami [140, 138], which includes post-Newtonian corrections to the acceleration up to the 3rd order.
As we can see, the torque time series presents several oscillating components at different harmonics of the inner binary’s orbital frequency. The amplitude of the oscillations typically exceeds the orbit-averaged value by a factor $\sim 5$. The two major oscillatory components can by attributed to inner binary’s and the outer binary’s orbital motions, respectively. A more detailed view of the various components is presented in the bottom panel of Fig. 8, in which we express the torque time series as a periodogram in the form of Eq. II.1. We observe that the torque spectrum is composed by several distinctive peaks, the most prominent of which occur at $\sim 1.8f_{\rm K}$ and $\sim 0.086f_{\rm K}$ respectively. The exact value and relation of these two peaks encodes information about the relative orbital parameters of the binary and the tertiary. In particular, resolving such torque peaks would yield information regarding the mass, distance and eccentricity of the latter, and allow to break the perturber mass-distance degeneracy that exists for secular de-phasing in triples. We note that this specific realisation of a triple system can be easily scaled to arbitrary masses and separation by approximating the potential of the tertiary as a tidal field:
@@
where $m_{3}$ is the tertiary’s mass and $R$ is the distance of the binary to the tertiary. Like other secular tidal effects, time varying oscillations will therefore be stronger when the perturber is massive and close-by. Such time-varying forces always accompany any other effect caused by the presence of a tertiary body.
@@ V.3 Massive black holes in circumbinary discs
The dynamics of gas-embedded comparable-mass systems are highly non-trivial and have primarily been studied numerically [for a recent review see 77]. A primary focus of such investigations has been the periodicity of the accretion rate onto the binary (which is not always dominated by the binary’s orbital frequency) and its applications to electromagnetic searches for compact massive binaries [e.g. 48, 57]. In [106] a spectral analysis of the torques in such systems was performed, but nearly all recent analyses of the gas forcing and resultant change in orbital elements have focused on the time-averaged, secular evolution of the orbital elements.
In the top panel of Fig. 9 we illustrate a raw time-series of the torques experienced by an equal-mass binary fixed on a circular orbit over 40 periods. We see that the resultant torques in a quasi-steady configuration are in fact highly variable, with features beyond those solely at the binary orbital frequency. Here the magnitude of the torque shows fluctuations by a factor of order $\sim 30$ with respect to its orbit-averaged value. The presented data is taken from a 2D, isothermal-hydrodynamics simulation performed with Sailfish [c.f. 143]. The setup is exactly that presented in [52] for a thin, co-planar disc of characteristic scale-height $h/r\approx\mathcal{M}^{-1}=0.1$ and global kinematic viscosity $\nu=\nu_{\rm d}a^{2}\Omega_{b}$ with $\nu_{\rm d}=10^{-3}$ (roughly equivalent to a turbulo-viscous parameter $\alpha=0.1$ at $r=a$; [121]).
The bottom panel of Fig. 9 shows a periodogram of the components of the torque. In comparison to three-body systems, hydro-dynamical torques are significantly more complex and spread over a wide range of harmonics. Therefore, we expect the $\delta$SNR of the resulting phase perturbations to be significantly larger than what individual torque component peaks may suggest (following Eq. 51). Nevertheless, we can identify several peaks at approximately $\sim 0.38f_{\rm K}$, $\sim 1.6f_{\rm{K}}$ and $\sim 3.9f_{\rm K}$. Therefore, a high-SNR observation of GWs from such a system would in principle allow one to reconstruct features in the flow of gas around in-spiralling black holes. We note that this should be the case for any type of gas-embedded binary, including stellar-mass binaries in accretion discs [e.g. 5, 79, 37, 107, 144, 43].
@@
where $\dot{M}$ is the accretion rate onto the binary. The parameter $q_{\rm d}$ can be roughly understood as the ratio of the disc’s scale mass and the binary total mass, which depends on the binary separation [compare with e.g. 131]. For realistic accretion discs this parameter will be small when sources enter the m-Hz band, of order $\sim 10^{-5}$ to few $\sim 10^{-1}$, depending on the total binary mass and the properties of the disc.
@@ V.4 EMRIs and IMRIs in accretion discs
Simulations exploring E/IMRI coalescence in accretion discs find that the gas exerts torques with fluctuating magnitude on the embedded BH, which roughly averages out to match the expectations of linear perturbation theory and planetary migration over many orbits [129, 64, 38]. In Fig. 10, we show a time-series of the instantaneous torque experienced by an IMRI embedded in a steady-state, iso-thermal and viscous accretion disc. The data are taken from un-smoothed results of the simulations detailed in [38], i.e. 2D hydro-dynamical Newtonian simulation performed with the code DISCO [51], in which an IMRI slowly in-spirals on a prescribed, quasi-circular orbit. The parameters for the simulation are a binary mass-ratio $q=10^{-3}$, a disc alpha viscosity parameter $\alpha=0.03$, and a uniform disc scale height $h/r\approx 0.03$. Here we additionally choose a conservative surface density scale for the disc of $0.1$ g/cm2 [compare with e.g., 38], which already suffices to produce very strong effects. Note that both the secular torques and the fluctuations scale linearly with the disc’s scale density.
@@ VI Conclusions
  • A complete data analysis pipeline that takes all GW parameters into account, including spins, various orientation angles and potentially the ulterior GW perturbations discussed in Sec. II.
@@
In the meantime, we suggest that the latter be included in future studies of EEs by means of Eqs. 27 and 28 (or any other extension), in addition to secular de-phasing prescriptions.
@@ Acknowledgements.
L.Z. acknowledges support from ERC Starting Grant No. 121817–BlackHoleMergs. A.A.T. acknowledges support from the European Union’s Horizon 2020 and Horizon Europe research and innovation programs under the Marie Skłodowska-Curie grant agreements No. 847523 and 101103134. D.J.D. and C.T. received support from the Danish Independent Research Fund through Sapere Aude Starting grant No. 121587. Z.H. acknowledges support from NSF grant AST-2006176 and NASA grants 80NSSC22K0822 and 80NSSC24K0440. AD acknowledges support from NSF grant AST-2319441.

### v3

@@ IV.1.1 Comparison with secular dephasing
The previous equation speaks to the high potential detectability of secular dephasing. It suggests that, indeed, cumulative shifts of order few $1/\text{SNR}$ may be detectable. However, this conclusion ignores the fact that secular dephasing, by its own nature of being slowly accumulated, can be degenerate with many binary parameters that act in a similar way. In particular, it is highly degenerate with the binary’s initial phase, along with any post-Newtonian modifications to the binary’s chirp behaviour with a similar magnitude or frequency scaling. A more conservative requirement for the detectability of secular dephasing is that it should exceed a full cycle, i.e. $2\pi$, while many works have also highlighted cases where several cycles of dephasing are necessary to perform parameter estimation [see, e.g., 130, 76]. Here we use $2\pi$ as a detectability criterion, finding that the threshold acceleration for secular de-phasing amounts to:
@@
where $N_{\rm{c}}$ is the number of cycles at the frequency $f$. Interestingly, this suggests that periodic perturbations may be the dominant EE for high SNR sources that complete comparatively fewer cycles in the sensitivity band of their respective detectors. As we will see in Sec. V, this may be realised in several astrophysical scenarios. The equation above also suggests that, in sources where both secular and periodic forces are present, the latter must be significantly higher in amplitude to be detected at the same $\delta$SNR. However a more precise statement on detectability requires a careful Bayesian analysis. Indeed, as we will see in IV.4, these naive expectations are often misleading.
@@ IV.2 Detectability for chirping LISA sources
When stepping beyond the monochromatic case, it is necessary to account for the full noise power spectral density of a detector. Here we produce some simple detectability criteria for the recently adopted, space based gravitational wave detector LISA [4, 34], but note that similar estimates apply to any similar milli-Hz GW instrument [e.g. TianQin 93].
@@
$\sim 8$ $\implies\text{1$\sigma$ uncertainty $<$ 0.5 for $\log_{10}(B^{\prime})$ and $\beta^{\prime}$}.$
@@ IV.4 Test case: Degeneracies in secular vs periodic dephasing
The test case showcased in Fig. 6 is representative of all $\delta$SNR larger than a few, where the widths of the Gaussian posteriors roughly scale as $1/\delta$SNR. This suggests that the additional parameters required to describe periodic de-phasing are completely orthogonal to the usual vacuum binary parameters. Crucially, they do not cause any visible degradation in the recovery of the vacuum waveform. Note that here we used a value of $n=1$ as a proxy for any perturbation on close to orbital timescales, still neglecting the secular effect of resonances. The results for the posteriors with diferent values of n roughly follow the expectations given by the $n^{-2}$ scaling of Eq. 29.
@@
We are aware that a complete validation of this claim would require a more systematic study of both secular and periodic dephasing, for many sources and across an array of different waveform models, which is beyond the scope of this work. Nevertheless, the results above do suggest that waveforms of the type presented in Eqs. 27 and II.2 represent a strong candidate to search for EE in GW signals without necessarily degrading the estimation of the vacuum parameters of the source.
Figure 7: Full showcase of the marginalised posteriors when attempting to recover a periodic (black) and a secular (red) dephasing. The injected periodic dephasing is characterised by $B^{\prime}=0.1$ and $\beta$ has a $\delta$SNR of approximately 40. the injected secular dephasing has the same amplitude and scaling, resulting in a $\delta$SNR of approximately 60. Nevertheless, the latter cannot be reconstructed confidently as it suffers from a degeneracy with the binary’s initial phase and spin parameter. The two MCMCs are performed with 32 parallel walkers for 50’000 steps, requiring approximately 90 minutes of computation on a single core. See Eqs. 44 for the meaning of the nine parameters.
@@ V.1 Generic multipole perturbations
We apply the formalism developed in this work to a very simple, heuristic description of multi-polar modifications to the gravitational potential. We imagine the vacuum gravitational field to be perturbed by a series of additional components, arising either from the presence of background fluid, field or from modifications to GR (see e.g. [37, 38] for a recent example). As the binary orbits in the modified potential, it will experience time varying forces with the following physical scaling:
@@
Extrapolating from the monochromatic results discussed in section IV.1 we conclude that by employing waveforms of the type showcased by Eqs. 27 and II.2, one should expect to be able to constrain completely general multipolar modifications to the gravitational potential (and therefore the spacetime metric), provided they exceed a certain typical magnitude:
@@ V.2 Dynamical and stellar triple systems
Triple systems of compact objects are a rich laboratory of gravitational dynamics, and are being studied in great detail in the context of the dynamical formation channel for GW sources [104, 17, 122, 88, 110, 111, 111, 8, 100, 7, 128, 117, 118, 140, 109, 87, 71, 119, 147, 145, 70, 21, 145]. It is known that a close-by tertiary massive body may induce several detectable imprints on GW emission from the inner binary. In this context, the most studied EEs are e.g. Doppler modulations that arise as a consequence of Römer delays or line-of-sight accelerations, as well as the secular influence of tidal fields throughout the inspiral [160, 20, 115, 73, 108, 28, 105, 154, 134, 53, 143, 132, 157, 81, 139, 121, 121], such as the secular drift caused by the von Zeipel-Kozai-Lidov mechanism [149, 79, 85]. Here we neglect secular effects, focusing instead on the instantaneous forces experienced by the inner binary components, as they orbit in the time-varying gravitational potential induced by the tertiary.
In the top panel of Fig. 8 we show the time series of the instantaneous torque experienced by an inner binary due to its tertiary companion. In this specific case, the inner binary has masses of 25 and 20 $M_{\odot}$, while the outer companion has a mass of 30 $M_{\odot}$. The inner binary has a semi-major axis of $2.5\times 10^{-4}$ AU, and emits GWs at 0.1 Hz. The outer binary has a semi-major axis of $1.5\times 10^{-3}$ AU, yielding a period ratio of approximately 11. The triple is coplanar, so that the von Zeipel-Kozai-Lidov mechanism does not take place. The simulations are performed with the regularized, direct $N$-body code tsunami [146, 144], which includes post-Newtonian corrections to the acceleration up to the 3.5 order.
As we can see, the torque time series presents several oscillating components at different harmonics of the inner binary’s orbital frequency. The amplitude of the oscillations typically exceeds the orbit-averaged value by a factor $\sim 5$. The two major oscillatory components can by attributed to inner binary’s and the outer binary’s orbital motions, respectively. A more detailed view of the various components is presented in the bottom panel of Fig. 8, in which we express the torque time series as a periodogram in the form of Eq. II.1. Note that this specific realisation of a triple system can be easily scaled to arbitrary masses and separation by approximating the potential of the tertiary as a tidal field:
@@
where $m_{3}$ is the tertiary’s mass and $R$ is the distance of the binary to the tertiary. We observe that the torque spectrum is composed by several distinctive peaks, the most prominent of which occur at $\sim 1.8f_{\rm K}$ and $\sim 0.086f_{\rm K}$ respectively. The exact value and relation of these two peaks encodes information about the relative orbital parameters of the binary and the tertiary. Like other secular tidal effects, time varying oscillations will therefore be stronger when the perturber is massive and close-by and are always accompanied to any other effect caused by the presence of a tertiary body. However, resolving such torque peaks would uniquely yield detailed information regarding the mass, distance and eccentricity of the latter, allowing to break the perturber mass-distance degeneracy. As an example, the low-frequency perturbation is a consequence of the tertiary’s orbit: its exact frequency scales as the ratio of periods $\propto(p_{0}/R)^{3/2}$, while its amplitude scales as $\propto Gm_{3}p_{0}/R^{3}$ and is also mediated by the tertiary’s eccentricity.
@@ V.3 Massive black holes in circumbinary discs
The dynamics of gas-embedded comparable-mass systems are highly non-trivial and have primarily been studied numerically [for a recent review see 82]. A primary focus of such investigations has been the periodicity of the accretion rate onto the binary (which is not always dominated by the binary’s orbital frequency) and its applications to electromagnetic searches for compact massive binaries [e.g. 51, 61]. In [112] a spectral analysis of the torques in such systems was performed, but nearly all recent analyses of the gas forcing and resultant change in orbital elements have focused on the time-averaged, secular evolution.
In the top panel of Fig. 9 we illustrate a raw time-series of the torques experienced by an equal-mass binary fixed on a circular orbit over 40 periods, which arise from both gravitational forces and gas accretion and vary quasi-periodically at multiple frequencies. We see that the resultant torques in a quasi-steady configuration are in fact highly variable, with features beyond those solely at the binary orbital frequency. Here the magnitude of the torque shows fluctuations by a factor of order $\sim 30$ with respect to its orbit-averaged value. The presented data is taken from a 2D, isothermal-hydrodynamics simulation performed with Sailfish [c.f. 150]. The setup is exactly that presented in [55] for a thin, co-planar disc of characteristic scale-height $h/r\approx\mathcal{M}^{-1}=0.1$ and global kinematic viscosity $\nu=\nu_{\rm d}a^{2}\Omega_{b}$ with $\nu_{\rm d}=10^{-3}$ (roughly equivalent to a turbulo-viscous parameter $\alpha=0.1$ at $r=a$; [127]).
The bottom panel of Fig. 9 shows a periodogram of the components of the torque. In comparison to three-body systems, hydro-dynamical torques are significantly more complex and spread over a wide range of harmonics. Therefore, we expect the $\delta$SNR of the resulting phase perturbations to be significantly larger than what individual torque component peaks may suggest (following Eq. 50). Nevertheless, we can identify several peaks at approximately $\sim 0.38f_{\rm K}$, $\sim 1.6f_{\rm{K}}$ and $\sim 3.9f_{\rm K}$. Therefore, a high-SNR observation of GWs from such a system would in principle allow one to reconstruct features in the flow of gas around in-spiralling black holes. We note that the numerical tests illustrated in section IV strongly suggest that periodic perturbations will be orthogonal the secular de-phasing arising from accretion or gravitational torques, which necessarily accompany them. The latter have been studied in detail in e.g. [66, 162] and [137]. The relative importance of periodic versus secular perturbations will strongly depend on the properties of the accretion disc. A joint detection of secular dephasing and periodic perturbations would be an unprecedented probe of the disc’s structure, and will be the focus of future work. We also note that our qualitative statements should hold for any type of gas-embedded binary, including stellar-mass binaries in accretion discs [e.g. 5, 84, 40, 113, 151, 46].
@@
where $\dot{M}$ is the accretion rate onto the binary. The parameter $q_{\rm d}$ can be roughly understood as the ratio of the disc’s scale mass and the binary total mass, which depends on the binary separation [compare with e.g. 137]. For realistic accretion discs this parameter will be small when sources enter the m-Hz band, of order $\sim 10^{-5}$ to few $\sim 10^{-1}$, depending on the total binary mass and the properties of the disc.
@@ V.4 EMRIs and IMRIs in accretion discs
Simulations exploring E/IMRI coalescence in accretion discs find that the gas exerts torques with fluctuating magnitude on the embedded BH, which roughly averages out to match the expectations of linear perturbation theory and planetary migration over many orbits [135, 68, 41]. In Fig. 10, we show a time-series of the instantaneous torque experienced by an IMRI embedded in a steady-state, iso-thermal and viscous accretion disc. The data are taken from un-smoothed results of the simulations detailed in [41], i.e. 2D hydro-dynamical Newtonian simulation performed with the code DISCO [54], in which an IMRI slowly in-spirals on a prescribed, quasi-circular orbit. The parameters for the simulation are a binary mass-ratio $q=10^{-3}$, a disc alpha viscosity parameter $\alpha=0.03$, and a uniform disc scale height $h/r\approx 0.03$. Here we additionally choose a conservative surface density scale for the disc of $0.1$ g/cm2 [compare with e.g., 41], which already suffices to produce very strong effects. Note that both the secular torques and the fluctuations scale linearly with the disc’s scale density.
@@ VI Conclusions
  • A complete data analysis pipeline that takes all GW parameters into account, including spins, various orientation angles and potentially the additional GW perturbations discussed in Sec. II.
@@
In the meantime, we suggest that the latter be included in future studies of EEs by means of Eqs. 27 and II.2 (or any other extension), in addition to secular de-phasing prescriptions.
@@ Acknowledgements.
L.Z. acknowledges support from ERC Starting Grant No. 121817–BlackHoleMergs. A.A.T. acknowledges support from the European Union’s Horizon 2020 and Horizon Europe research and innovation programs under the Marie Skłodowska-Curie grant agreements No. 847523 and 101103134. D.J.D. and C.T. received support from the Danish Independent Research Fund through Sapere Aude Starting grant No. 121587. Z.H. acknowledges support from NSF grant AST-2006176 and NASA grants 80NSSC22K0822 and 80NSSC24K0440. AD acknowledges support from NSF grant AST-2319441. The authors thank the referee for many useful comments.